### v1

# CaloQVAE : Simulating high-energy particle-calorimeter interactions using hybrid quantum-classical generative models

Abstract
The Large Hadron Collider’s high luminosity era presents major computational challenges in the analysis of collision events. Large amounts of Monte Carlo (MC) simulation will be required to constrain the statistical uncertainties of the simulated datasets below these of the experimental data. Modelling of high-energy particles propagating through the calorimeter section of the detector is the most computationally intensive MC simulation task. We introduce a technique combining recent advancements in generative models and quantum annealing for fast and efficient simulation of high-energy particle-calorimeter interactions.
preprint: APS/123-QED
[…]
The significant increase in the rate of data-taking presents a computational challenge in the generation of sufficient high-quality simulated datasets, which are typically used to both calibrate the detectors and evaluate compatibility of the data with different physical hypotheses. These simulated datasets are produced with first-principles particle simulation software in the GEANT4 framework [3], though less-accurate, parameterized models of the detectors, often called “fast simulation”, are also used [4]. The increased rate of data-taking and precision required to exploit the high-statistics HL-LHC dataset demands both that the computational time to produce simulation decrease, and that the quality of the simulation remains as close as possible to the first-principles simulation [5].
A significant portion of the computational burden of the simulation lies in the interaction of particles with the calorimeter subsystems. Calorimeters are detectors that measure particle energies via the production of showers and the subsequent measurement of those showers as they traverse the active material of the detector. The ATLAS and CMS experiments both contain “Electromagnetic” calorimeters dedicated to the measurement of photons and electrons, and “Hadronic” calorimeters dedicated to measuring hadrons (typically $\pi^{\pm}$, protons, and neutrons). The accurate simulation of these complicated particle showers is critical to enable the highest quality measurements, but simulating each shower from first principles is very time consuming. This has motivated in recent years several approaches to reproduce these showers via machine learning based methods (see below).
In this paper, we demonstrate the use of a quantum-assisted machine learning model for the generation of particle showers at LHC-like detectors. We deploy a restricted Boltzmann machine (RBM) to encode a rich description of particle showers in detectors, and use quantum annealing to avoid expensive Markov Chain evaluations of the shower model. The technique shows promise in both increasing the fidelity of fast simulation, and reducing the computational requirements via a new computing paradigm.
Recently, significant advancements have been made in various machine learning domains related to generative models. In the development of variational autoencoders (VAEs) and their extensions, a series of breakthroughs have shaped the field. The introduction of auto-encoding variational Bayes in 2014 brought about a stochastic variational inference algorithm that enabled handling of intractable posterior distributions and large datasets [6]. Expanding upon VAEs, discrete variational autoencoders (DVAEs) emerged in 2016, integrating discrete and continuous latent variables through smoothing transformations and hierarchical structures, thus enhancing their representational power [7]. Building upon these foundations, overlapping transformations were introduced in 2018 to refine the training of discrete variables in DVAEs, resulting in state-of-the-art performance [8]. Around the same time, GumBolt was proposed as a model for discrete latent variables with Boltzmann machine priors, leveraging the Gumbel trick to capture complex distributions over discrete variables [9]. Collectively, these advancements significantly contribute to the field of VAEs and their extensions, addressing challenges related to intractable posteriors, incorporating discrete variables, and leveraging different priors for enhanced generative modelling.
The development of these generative models provides inspiration to collider physics in the context of calorimeter simulation. In 2018, novel techniques based on generative adversarial networks (GANs) were developed to address the need for fast simulation of electromagnetic showers in calorimeters [10]. In 2021, the advancements have introduced innovative frameworks like CaloFlow based on normalizing flows, offering fast and accurate simulations with high fidelity and stability [11]. Based on the successes of these studies, we explore whether quantum computing, particularly the D-Wave processor, can contribute to generative model strategies for collider physics simulation.
In 2015, a method utilising quantum annealing for training deep neural networks was proposed with a description on how to map RBMs, the basic building blocks of deep belief networks (DBNs), onto the D-Wave Two hardware [12]. Building on this idea, a subsequent work explored quantum annealers’ ability to sample from Boltzmann distributions, overcoming several challenges including effective temperature estimation and restricted quantum hardware architecture, to propose an algorithm to enhance RBM learning with quantum hardware [13]. In 2017, a study tested the freeze-out conjecture in quantum annealers, revealing most instances do not exhibit a well-defined freeze-out point or an effective temperature due to quantum fluctuations and noise [14]. The subsequent development of the quantum variational autoencoder (QVAE) showcased the feasibility of utilising a quantum Boltzmann machine as a prior for latent generative process in a variational autoencoder [15, 16]. The study of QVAE also highlighted the potential of large latent-space quantum Boltzmann machines in achieving quantum advantage for deep generative models. These innovations serve to motivate a new research direction: employing quantum generative models for simulating High-Energy Physics (HEP) detectors. In 2021, a hybrid quantum-classical qGAN showed promise for accelerating HEP detector simulations using quantum generator circuits and classical discriminator neural networks [17]. In a very recent development, CaloDVAE introduced a discrete variational autoencoder with an RBM prior for fast calorimeter shower simulation, yielding promising results in generating realistic and diverse samples. The possibility of using quantum annealing processors as sampling devices for the RBM prior was also discussed, which represents a crucial initial step towards applying QVAEs for calorimeter shower simulation and exploring quantum annealing processors in HEP research [18].
Variational Autoencoders (VAEs) are a class of latent variable generative models, $p_{\theta}(\mathbf{x},\mathbf{z})$, that maximize an evidence lower bound (ELBO) to the true log-likelihood:
$\text{log }p_{\theta}(\mathbf{x})\geq\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}[\text{log }p_{\theta}(\mathbf{x}|\mathbf{z})]-\text{KL}[q_{\phi}(\mathbf{z}|\mathbf{x})||p_{\theta}(\mathbf{z})]$ (1)
$\text{KL}[Q||P]$ is the Kullback-Liebler divergence between two probability distributions, $Q$ and $P$ while $\mathbb{E}_{p}$ denotes an expectation value over the distribution $p$. The variables $\mathbf{x}$ and $\mathbf{z}$ represent data (in this case a vector of calorimeter cell energies) and latent variable respectively. The approximating posterior, $q_{\phi}(\mathbf{z}|\mathbf{x})$ and the generative, $p_{\theta}(\mathbf{x}|\mathbf{z})$ distributions are often parameterized using neural networks. In the original VAE framework [6, 19], $q_{\phi}(\mathbf{z}|\mathbf{x})$ and $p_{\theta}(\mathbf{z})$ are assumed to be factorized Gaussian distributions, $\mathcal{N}(\bm{\mu}_{\phi}(\mathbf{x}),\bm{\Sigma}^{2}_{\phi}(\mathbf{x}))$ and $\mathcal{N}(\mathbf{0},\mathbf{I})$ respectively.
CaloDVAE [7, 8] is a hierarchical discrete VAE which extends the VAE framework by introducing discrete latent variables $\mathbf{z}_{i}\in\{0,1\}$ in the latent space. Both the encoder, $q_{\phi}(\mathbf{z}|\mathbf{x},\mathbf{e})$, and decoder, $p_{\theta}(\mathbf{x}|\mathbf{z},\mathbf{e})$, are conditioned on energy [20, 18] and are modelled by fully connected neural networks. The approximate posterior of the VAE has a hierarchical structure $q_{\phi}(\mathbf{z}|\mathbf{x},\mathbf{e})=\prod_{i}q_{\phi_{i}}(\mathbf{z}_{i}|\mathbf{z}_{j<i},\mathbf{x},\mathbf{e})$ with $N$ latent groups and the latent distribution of the VAE is modelled by an RBM where $\mathbf{z}=[\mathbf{z}_{1},...,\mathbf{z}_{N}]$ are partitioned into $2$ equal subsets to make the $2$ sides of the RBM [18]. The RBM is modeled by probability distribution $p_{RBM}=e^{\mathbf{a}_{l}^{T}\mathbf{z}_{l}+\mathbf{a}_{r}^{T}\mathbf{z}_{r}+\mathbf{z}_{l}^{T}\mathbf{W}\mathbf{z}_{r}}/Z$ where $Z$ is the partition function and $(\mathbf{a}_{l},\mathbf{a}_{r},\mathbf{W})$ are the RBM parameters which are trained along with the VAE parameters $(\theta,\phi)$.
The VAE and RBM parameters are jointly trained using the standard VAE objective and a binary cross-entropy (BCE) loss to enhance the generative capability of the model. Firstly, binary labels are created to isolate pixel positions with non-zero energy values in the energy images. Subsequently, a BCE loss is computed between the binary label of the preprocessed input data and the binary label of the reconstructed data to incentivize the VAE to faithfully reproduce the distribution of non-zero energy pixels [18].
Due to non-differentiability of discrete values we replace discrete values with continuous variables following the Gumbel trick to allow backpropagation during training [9]. Therefore the latent discrete variables $\mathbf{z}$ are replaced with continuous proxy variables $\zeta$ with the Gumbel trick. Similarly we also apply Gumbel trick on reconstructed labels to mask the reconstructed energies. Finally in inference the continuous variables are replaced with the discrete variables [18].
In CaloQVAE, we propose a novel calorimeter simulation framework based on the quantum variational autoencoder generative model. In CaloDVAE, block Gibbs sampling is used to generate latent RBM samples from a trained model. In this work we show that we can generate latent space samples using the D-Wave 2000Q annealer. To generate samples using the quantum processing unit (QPU), we first need to embed the RBM on the quantum hardware. However, the QPU architecture is not fully connected and instead follows a Chimera graph topology, making it impossible to embed a fully connected RBM in the QPU [12]. We therefore create a masking function to remove selected edges from an RBM to produce a Chimera-constrained RBM. We train the model classically analogously to the CaloDVAE training by replacing the RBM of CaloDVAE with a Chimera-constrained RBM.
[…]
Figure 1: Histogram showing the probability of obtaining sample $s$ with Ising Energy computed by $E(s)=s^{T}Js+h^{T}s$ in dimensionless units for a trained CaloQVAE model. The energy distribution of QPU samples (solid line) are very close to the Classical samples (shaded) after we account for the $\beta_{eff}^{*}$ factor.
For training and performance evaluation we use the public CaloGAN dataset, which provides GEANT4-simulated datasets with 100,000 events for $\pi^{+}$ and $e^{+}$ incident particles [10]. The data is split into 80%, 10%, and 10% subsets for training, validation, and testing, respectively. After conducting a hyperparameter scan [18], we obtain the best models for each particle type. Utilizing classical blocked Gibbs sampling and QPU sampling methods respectively, we generate new samples from the latent space and passed them through the decoder to obtain synthetic shower images. These images are then compared with GEANT4 test data to assess their performance. Classification can be used as a useful indicator of the performance of generative models. A good classifier should be able to understand and capture the important patterns in the data because it needs to learn discriminating features that distinguish different classes or categories accurately. We perform binary classification of $e^{+}$ vs $\pi^{+}$ and $e^{+}$ vs $\gamma$ using a neural network with 6 fully connected layers with dropout, following the method of [10]. All synthetic data is generated from the QPU using separate models trained separately on $e^{+},\pi^{+}$ and $\gamma$ samples. Our accuracy metrics (Table 1) are very close to the results obtained in [10]. The similar accuracy metrics illustrate that our CaloQVAE model is generating synthetic data that captures important patterns present in the true data, which indicates that the synthetic images generated from the QPU are successful at representing the underlying structure and characteristics of the true data. The fact that the classifier trained on synthetic $e^{+}$ vs $\pi^{+}$ data can perform well when tested on the GEANT4 data, and vice versa, suggests that the features learned by the CaloQVAE model are transferable. This means that the CaloQVAE model has successfully encoded the relevant information from the true data into its latent space representation, which allows the generated data to be informative and useful for training a classifier. An issue that requires attention is the unexpected high performance of the $e^{+}$ vs $\gamma$ classifier, which is trained with synthetic data and tested on a separate set of synthetic data. Similar over-performance has also been observed in [10]. To address this matter, we first check the datasets and discovered that the mean energy distributions on 504 calorimeter cells are similar for both the $e^{+}$ and $\gamma$ synthetic datasets. We then explored potential solutions such as masking out the low-energy pixels and unifying the normalization of the datasets for the two types of particles. However, despite these efforts, we are unable to effectively resolve the issue of over-performance. This raises the possibility that our CaloQVAE model may generate certain non-physical features, which allow the classifier to distinguish between the $e^{+}$ and $\gamma$ samples. Further research is warranted to explore potential ways of addressing this challenge.
$e^{+}$ vs $\pi^{+}$ $e^{+}$ vs $\gamma$
Test Test
Geant4 Synthetic Geant4 Synthetic
Train Geant4 99.8 99.9 66.8 74.1
Synthetic 94.3 100.0 53.7 99.9
Table 1: Binary classification accuracy for $e^{+}$ vs $\pi^{+}$ and $e^{+}$ vs $\gamma$ where we train and test both on the Geant4 data and synthetic data produced from QPU samples
We qualitatively analyze one-dimensional histograms of the typical shower shape variables. Included in this publication are the fractions of the energy recorded in the first calorimeter layer and the fraction of first layer with any energy deposited. The latter is also called the sparsity. As shown in Fig 2 there is a close alignment between the results from GEANT4 data and generative models, which signifies that our generative model effectively captures the underlying patterns and distributions of the data. The consistent distribution patterns highlight the generative model’s fidelity in representing the essential features and correlations present in the training dataset. Moreover, the similarity between the performance of QVAE and DVAE models indicates the quality of samples from the quantum device is at par with modern Monte Carlo methods.
Figure 2: Histogram of Shower Shape Variables.
To evaluate the performance of model’s energy conditioning needed for practical deployment, we sample from the models requesting a specific range of true energy (45-60 GeV) of the incident particle and histogram the total observed energy in the positron and pion clusters. We sample with both classical approach (DVAE) and quantum approach (QVAE) and comparing the observed energy with true energy from the same range in the GEANT4 test dataset. As shown in Fig 3, for the positron cluster, the conditioning response of DVAE or QVAE can form a flat plateau with good accuracy in reproducing the defined selection of the input energy range. The pion cluster has a broadened response that is due to the nature of the unconstrained charged pion shower in the electromagnetic calorimeter. The histograms of the generative models for both two clusters are able to mirror the characteristics observed in the true data. This compelling agreement underscores the VAE’s ability to capture and faithfully reproduce the intricate patterns and correlations inherent in the first-principles simulation, thereby affirming its efficacy as a powerful tool for energy pattern generation.
Figure 3: Observed energy spectra for synthetic CaloDVAE compared with GEANT4 test data.
The monitoring results of the DVAE sampling rate show efficient performance on our NVIDIA GeForce RTX 2080 Ti with 11264 MB memory. The processing time per sample for latent space sampling takes the most substantial portion ( $\sim$ 0.5 $\mathrm{ms}$), followed by data loading to the GPU ( $\sim$ 10 $\mathrm{ns}$), and passing through the decoder ($\sim$ 1 $\mathrm{\mu s}$). The DVAE model exhibits similar sampling rate with the CaloGAN results [10] up to an order of magnitude. On the other hand, the quantum sampler achieves slightly faster sampling rate at 0.4 $\mathrm{ms}$ per sample, but impressively, the core QPU annealing rate is only 20 $\mathrm{\mu s}$ per sample, hinting at the raw speed of quantum processing. The extra QPU access time is mainly used for tasks like readout and handling delays, suggesting that with optimized engineering, quantum sampling could vastly outpace classical methods in efficiency for generative modeling, making it well-suited for real-time and large-scale applications. We expect to apply the QVAE model to more complex objects, jets and events in the future, for which annealing time is expected to remain at the current level.
This work is the first demonstration of the application of a quantum annealing device to the computationally expensive simulation of particle showers at the Large Hadron Collider. We have shown that it is possible to utilize the Quantum Processing Unit (D-Wave Chimera 2000Q) to generate RBM samples which can be used to generate particle showers. The resulting generated simulation is able to largely match the properties of the training dataset, and match the performance of other generative models. The raw QPU annealing time per sample is measured to be 20 $\mathrm{\mu s}$, which is 20 times faster than generating samples via GPU. While further work remains in evaluating on more realistic simulation and a wider range of physics topologies, the initial results are demonstrate CaloQVAE to be a promising application of quantum computing to open research questions in fundamental physics. The experimental results and code are available online [27].
Acknowledgements.
The authors would like to thank Drs. Walter Vinci, Mohammad Amin, Geoffrey Fox, the members of the DeGeSim Collaboration as well as the members of the Perimeter Institute Quantum Intelligence Lab for many enlightening discussions. We thank Dr O. Di Matteo for early project conceptualization and material support for A. Abhishek during the final stages of the research. This work is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), National Research Council (NRC) and the Canadian Institute for Advanced Research (CIFAR) AI chair program. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Economic Development, Job Creation and Trade. We gratefully acknowledge funding from the NRC via Agreement AQC-002, NSERC, including via grants SAPPJ-2020-00032, SAPPJ-2022-00020, and RGPIN-2022-04609 This research was enabled in part by computational support provided by the Shared Hierarchical Academic Research Computing Network (SHARCNET) and the Digital Research Alliance of Canada. S.H. and H.J. contributed equally to this work and are listed in alphabetical order.